# 1st Pranav Kulkarni    2nd Leo Sünkel    3rd Michael Kölle

### v2

@@ Abstract
Efficient entanglement distribution is the cornerstone of the Quantum Internet. However, physical link parameters such as photon loss, memory coherence time, and gate error rates fluctuate dynamically, rendering static purification strategies suboptimal. In this paper, we propose an Adaptive Purification Controller (APC) that autonomously optimizes the entanglement distillation sequence to maximize the ”goodput,” the rate of delivered pairs meeting a strict fidelity threshold. By treating protocol selection as a resource allocation problem, the APC dynamically switches between purification depths and protocol families (e.g., BBPSSW vs. DEJMPS) to navigate the trade-off between generation rate and state quality. Using a dynamic programming planner with Pareto pruning, simulation results demonstrate that our approach eliminates the ”fidelity cliffs” inherent in static protocols and prevents resource wastage in high-noise regimes. Furthermore, we extend the controller to heterogeneous scenarios, demonstrating robustness for both multipartite GHZ state generation and continuous variable systems using effective noiseless linear amplification models. We benchmark its computational overhead, confirming real-time feasibility with decision latencies in the millisecond range per link.
@@ I Introduction
Long-distance quantum communication and distributed quantum computing rely on the distribution of high fidelity entanglement over noisy channels with finite coherence times. Physical channels and devices introduce loss, dephasing, and gate errors, so raw Bell pairs or multipartite entangled states must be transformed into fewer but more entangled copies using local operations and classical communication (LOCC) [1, 2]. In repeater architectures, entanglement generation, purification, and swapping are interleaved along multi-hop paths, and the timing of these operations interacts non-trivially with channel round-trip times, detector efficiencies, and memory decoherence.
Several quantum network simulators have emerged to study these effects, including NetSquid [3, 4, 5], SimQN [6, 7], SeQUeNCe [8, 9], and QuNetSim [10]. Many provide circuit-level or fidelity-level physical models for entanglement generation and swapping, and some include single-link purification protocols or link-layer purification selection modules. However, purification is often configured in a static way, such as a fixed number of BBPSSW rounds per link, or treated as a purely local primitive without a global optimization over routed paths or over different protocol families (BBPSSW versus DEJMPS versus higher-order recurrence or hashing). Recent work on purification protocol selection integrates device-level noise models and density matrix simulation in a link-layer module [11], but does not couple this to a path-level planner that co-optimizes purification, swapping, and timing.
This paper introduces an Adaptive Purification Controller (APC), a module in the KOSMOS (Koordinations-, Optimierungs- und Simulations-Software für Modulare Quanten-Systeme in 6G Netzwerken) quantum network simulator that leverages the low latency of future 6G Quantum Technology. The APC sits between routing and the quantum execution driver. Given a routed path and a target end-to-end fidelity, the APC:
  • models each link using a fidelity-based Werner or Bell diagonal description and closed-form recurrence primitives,
@@ II-A Entanglement Purification Protocols
Entanglement purification aims to transform many copies of a noisy entangled state into fewer copies of a more entangled state using LOCC [1]. For bipartite qubit systems, two widely used families of recurrence protocols are:
@@
  • the DEJMPS protocol of Deutsch, Ekert, Jozsa, Macchiavello, Popescu and Sanpera [2], which operates on Bell diagonal states and can handle non-Werner noise more robustly.
@@
$\rho_{W}(F)=F|\Phi^{+}\rangle\langle\Phi^{+}|+\frac{1-F}{3}\sum_{k\in\{\Phi^{-},\Psi^{+},\Psi^{-}\}}|k\rangle\langle k|,$ (1)
@@
DEJMPS operates on Bell diagonal states of the form
@@
For a single DEJMPS round, the coefficients update according to [2]
$p_{\text{succ}}^{\text{DEJ}}(\bm{\lambda})$ $=(\lambda_{1}+\lambda_{2})^{2}+(\lambda_{3}+\lambda_{4})^{2},$ (5)
$\lambda_{1}^{\prime}$ $=\frac{\lambda_{1}^{2}+\lambda_{2}^{2}}{p_{\text{succ}}^{\text{DEJ}}},\quad\lambda_{2}^{\prime}=\frac{2\lambda_{1}\lambda_{2}}{p_{\text{succ}}^{\text{DEJ}}},$ (6)
$\lambda_{3}^{\prime}$ $=\frac{\lambda_{3}^{2}+\lambda_{4}^{2}}{p_{\text{succ}}^{\text{DEJ}}},\quad\lambda_{4}^{\prime}=\frac{2\lambda_{3}\lambda_{4}}{p_{\text{succ}}^{\text{DEJ}}}.$ (7)
@@
In the KOSMOS implementation, Bell-diagonal states are stored as a tuple $\mathrm{BellDiagonal}(a,b,c,d)$ ordered as $(\Phi^{+},\Psi^{+},\Psi^{-},\Phi^{-})$. Therefore, when relating (4)–(5) to code, the mapping is $(a,b,c,d)=(\lambda_{1},\lambda_{3},\lambda_{4},\lambda_{2})$. This is a basis permutation only (local Pauli-frame relabeling) and is equivalent to the standard DEJMPS recurrence.
@@ II-C Entanglement Swapping and Fidelity-based Models
Entanglement swapping transforms entangled pairs $(A,B)$ and $(C,D)$ into an entangled pair $(A,D)$ via a Bell state measurement on $(B,C)$ [20]. Fidelity-based entanglement models used in high-level simulators such as SimQN represent each link by a Werner parameter $w$ related to fidelity by
@@ III Related Work
SimQN [6, 7] offers both state-based and fidelity-based physical models, including an entanglement model with fidelity tracking and analytic operations for entanglement generation, swapping, and simple distillation. Its entanglement model provides closed-form recurrence maps for Bennett-style distillation on Werner and mixed Bell diagonal states, exponential decoherence of the Werner parameter in memory and over channels, and analytic swapping rules. The model is designed for speed and scalability, but it does not include a built-in dynamic planner for per-hop protocol choice or multipartite GHZ and CV distillation.
SeQUeNCe [8, 9] is an event-driven quantum network simulator that includes an entanglement management module with BBPSSW-style purification and entanglement swapping. Purification is available as a protocol component (BBPSSWProtocol and BBPSSW_BDS), and tutorials demonstrate fixed round schemes on short repeater chains. The Bell diagonal variant can be configured to behave as BBPSSW (with twirling) or DEJMPS (without twirling). Twirling applies random local Pauli rotations to symmetrize the state, effectively erasing specific error biases to yield a standard Werner state characterized solely by fidelity. SeQUeNCe exposes a method that computes the updated fidelity for a given input fidelity based on these configurations. However, SeQUeNCe does not provide a global path-level planner that jointly optimizes per-hop method, number of rounds, and end-to-end latency or resource cost.
@@
Beyond simulators, recent work on network-level purification and scheduling has studied how depolarizing noise, gate errors, storage times, and routing policies affect performance on complex topologies [22]. These works typically consider fixed purification protocols and focus on where to purify and how often. A complementary line of work optimizes repeater chain policies or cutoff thresholds using Markov decision processes or dynamic programming, but often without explicit purification choices. In parallel, Acharya et al. [11] proposed a purification protocol selection module for bipartite links that uses density matrix simulations under realistic errors to build heuristics for choosing BBPSSW, DEJMPS, or other protocols at the link layer. Their module optimizes over a local configuration space per hardware instance but does not combine these choices into a path-level planner that accounts for swapping and end-to-end objectives.
@@ IV-B1 Static Recurrence
On each link $\ell$, APC represents the current pair by either a scalar fidelity $F$ and Werner parameter $w$ or by a Bell diagonal vector $\bm{\lambda}$, depending on the configured physical model. Conversion between fidelity and Werner parameter is given by Eq. 13.
@@
for Werner inputs using (2)–(3), and, where configured, Bell diagonal updates $(\bm{\lambda},p_{\text{succ}})\mapsto(\bm{\lambda}^{\prime},p_{\text{succ}}^{\prime})$ using the DEJMPS relations. For $r$ rounds on link $\ell$ under protocol $P$, APC computes the sequence $F^{(j+1)}=F_{P}^{\prime}(F^{(j)})$ and the multi-round success probability $P_{\text{succ}}^{(r)}$ from (9). The expected raw entanglement cost per surviving pair after $r$ rounds is $C_{\text{pairs}}^{(r)}$ in (10).
@@ IV-B2 Local Operations and Memory Noise
$\mathcal{E}_{2}(\rho)=(1-p_{2})\rho+p_{2}\frac{\mathbb{I}_{4}}{4},$ (17)
with error probability $p_{2}$ and depolarizing strength $\lambda_{2q}\in[0,1]$. In KOSMOS, the user-facing parameters $p_{1}$, $p_{2}$, and $p_{\mathrm{meas}}$ denote Pauli-twirled per-operation error probabilities (single-qubit, two-qubit, and local measurement bit-flip, respectively). The backend converts these primitive error rates into an effective two-qubit depolarizing strength via
@@
and composes independent depolarizing channels by multiplying reliabilities, $1-\lambda_{\mathrm{tot}}=\prod_{i}(1-\lambda_{i})^{n_{i}}$.
Assuming $n_{1}$ single-qubit gates, $n_{2}$ two-qubit gates, and $n_{m}$ measurements per recurrence round, the per-round reliability factor is
$r_{\text{round}}=r_{1}^{n_{1}}r_{2}^{n_{2}}r_{m}^{n_{m}},\qquad$ (19)
$r_{1}=1-\lambda_{1q},\;\;r_{2}=1-\lambda_{2q},\;\;r_{m}=1-\lambda_{\mathrm{meas}}.$ (20)
@@
    • Werner tracking: $w^{\prime}\leftarrow(1-\lambda_{\mathrm{tot}})w^{\prime}$ and map back using (15),
@@
APC uses (21) between recurrence rounds and before swap operations. The wait time $t_{\text{wait}}$ includes classical round-trip delays and scheduling constraints and is provided by the planner based on the chosen objective and time model.
@@ IV-C Multipartite GHZ Stabilizer Distillation
APC includes a GHZ post stage that can be invoked after bipartite arms from a hub to multiple leaf nodes have been purified. The current implementation focuses on GHZ states constructed from Bell pairs meeting at a hub, guided by stabilizer-based protocols [13, 14, 15].
A noisy $N$-qubit GHZ state is modeled as
$\rho_{\text{GHZ}}=F_{\text{GHZ}}|\text{GHZ}_{N}\rangle\langle\text{GHZ}_{N}|+(1-F_{\text{GHZ}})\sigma_{\perp},$ (23)
@@
Instead of simulating the full stabilizer circuit, the backend uses an effective update rule consistent with the success/failure structure of graph state purification. In each GHZ distillation pass, a set of $m$ stabilizers is conceptually measured using ancillas of fidelity $F_{\text{anc}}$ and measurements with error probability $p_{\text{meas}}^{\text{GHZ}}$. Let $p_{\text{good}}=F_{\text{GHZ}}$ be the probability that the state lies in the GHZ subspace and $p_{\text{bad}}=1-p_{\text{good}}$ the probability of being outside it. The per check acceptance probabilities are approximated by
$q_{\text{good}}\approx F_{\text{anc}}(1-p_{\text{meas}}^{\text{GHZ}}),\qquad q_{\text{bad}}\approx q_{\min}\leq\frac{1}{2},$ (24)
where $q_{\min}$ is a small floor imposed for numerical stability. Over $m$ checks, the acceptance probabilities are $q_{\text{good}}^{m}$ and $q_{\text{bad}}^{m}$. The overall acceptance probability and updated GHZ fidelity are then
$p_{\text{succ}}^{\text{GHZ}}$ $\approx p_{\text{good}}q_{\text{good}}^{m}+p_{\text{bad}}q_{\text{bad}}^{m},$ (25)
$F_{\text{GHZ,out}}$ $=\frac{p_{\text{good}}q_{\text{good}}^{m}}{p_{\text{succ}}^{\text{GHZ}}}.$ (26)
The EPR cost is increased by $m$ ancilla pairs per pass, and the time is increased by $m$ rounds of operations and classical delays. This effective model yields qualitatively similar behavior to explicit stabilizer simulations while keeping APC within a fidelity based framework.

### v3

@@ Abstract
Efficient entanglement distribution is a cornerstone of the Quantum Internet. However, physical link parameters such as photon loss, memory coherence time, and gate error rates fluctuate dynamically, rendering static purification strategies suboptimal. In this paper, we propose an Adaptive Purification Controller (APC) that automatically optimizes the entanglement distillation sequence to maximize the goodput, i.e., the rate of delivered pairs meeting a strict fidelity threshold. By treating protocol selection as a resource allocation problem, the APC dynamically switches between purification depths and protocols (BBPSSW vs. DEJMPS) to navigate the trade-off between generation rate and state quality. Using a dynamic programming planner with Pareto pruning, simulation results show that our approach mitigates the ”fidelity cliffs” inherent in static protocols and reduces resource wastage in high-noise regimes. Furthermore, we extend the controller to heterogeneous scenarios, and evaluate it for both multipartite GHZ state generation and continuous-variable systems using effective noiseless linear amplification models. We benchmark its computational overhead, showing decision latencies in the millisecond range per link in our implementation.
@@ I Introduction
Long-distance quantum communication and distributed quantum computing rely on the distribution of high-fidelity entanglement over noisy channels with finite coherence times. Physical channels and devices introduce loss, dephasing, and gate errors, so raw Bell pairs or multipartite entangled states must be transformed into fewer but more entangled copies using local operations and classical communication (LOCC) [3, 7]. In repeater architectures, entanglement generation, purification, and swapping are interleaved along multi-hop paths, and the timing of these operations interacts non-trivially with channel round-trip times, detector efficiencies, and memory decoherence.
Several quantum network simulators have emerged to study these effects, including NetSquid [5], SimQN [4], SeQUeNCe [2], and QuNetSim [9]. Many provide circuit-level or fidelity-level physical models for entanglement generation and swapping, and some include single-link purification protocols or link-layer purification selection modules. However, purification is often configured in a static way, such as a fixed number of BBPSSW rounds per link, or treated as a purely local primitive without a global optimization over routed paths or over different protocol families (BBPSSW versus DEJMPS versus higher-order recurrence or hashing). Recent work on purification protocol selection integrates device-level noise models and density matrix simulation in a link-layer module [17], but does not couple this to a path-level planner that co-optimizes purification, swapping, and timing.
This paper introduces an Adaptive Purification Controller (APC), a module in the KOSMOS (Koordinations-, Optimierungs- und Simulations-Software für Modulare Quanten-Systeme in 6G Netzwerken) quantum network simulator that leverages the low latency expected of future 6G quantum technology. The APC sits between routing and the quantum execution driver. Given a routed path and a target end-to-end fidelity, the APC:
  • models each link using a fidelity-based Werner or Bell-diagonal description and closed-form recurrence primitives,
@@ II-A Entanglement Purification Protocols
Entanglement purification aims to transform many copies of a noisy entangled state into fewer copies of a more entangled state using local operations and classical communication (LOCC). For bipartite qubit systems, two widely used families of recurrence protocols are:
@@
  • the DEJMPS protocol of Deutsch, Ekert, Jozsa, Macchiavello, Popescu and Sanpera [7], which operates on Bell-diagonal states and can handle non-Werner noise more robustly.
@@
$\rho_{W}(F)=F|\Phi^{+}\rangle\langle\Phi^{+}|+\frac{1-F}{3}\sum_{k\in\{\Phi^{-},\Psi^{+},\Psi^{-}\}}|k\rangle\langle k|$ (1)
@@
DEJMPS operates on Bell-diagonal states of the form
@@
For a single DEJMPS round, the coefficients update according to
$p_{\text{succ}}^{\text{DEJ}}(\bm{\lambda})$ $=(\lambda_{1}+\lambda_{4})^{2}+(\lambda_{2}+\lambda_{3})^{2},$ (5)
$\lambda_{1}^{\prime}$ $=\frac{\lambda_{1}^{2}+\lambda_{4}^{2}}{p_{\text{succ}}^{\text{DEJ}}},\quad\lambda_{2}^{\prime}=\frac{\lambda_{2}^{2}+\lambda_{3}^{2}}{p_{\text{succ}}^{\text{DEJ}}},$ (6)
$\lambda_{3}^{\prime}$ $=\frac{2\lambda_{2}\lambda_{3}}{p_{\text{succ}}^{\text{DEJ}}},\quad\lambda_{4}^{\prime}=\frac{2\lambda_{1}\lambda_{4}}{p_{\text{succ}}^{\text{DEJ}}}.$ (7)
@@
In the KOSMOS implementation, Bell-diagonal states are stored as a tuple $\mathrm{BellDiagonal}(a,b,c,d)$ ordered as $(\Phi^{+},\Psi^{+},\Psi^{-},\Phi^{-})$. Therefore, when relating (4)–(5) to code, the mapping is $(a,b,c,d)=(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})$.
Implementation note (local basis choice). The DEJMPS primitive used by APC assumes this Bell-state ordering (equivalently, a fixed choice of local basis). It does not apply any additional coefficient sorting/permutation step before evaluating (5).
@@ II-C Entanglement Swapping and Fidelity-based Models
Entanglement swapping transforms entangled pairs $(A,B)$ and $(C,D)$ into an entangled pair $(A,D)$ via a Bell state measurement on $(B,C)$ [20]. Fidelity-based entanglement models used in high-level simulators such as SimQN[4] represent each link by a Werner parameter $w$ related to fidelity by
@@ III Related Work
SimQN [4] offers both state-based and fidelity-based physical models, including an entanglement model with fidelity tracking and analytic operations for entanglement generation, swapping, and simple distillation. Its entanglement model provides closed-form recurrence maps for Bennett-style distillation on Werner and mixed Bell diagonal states, exponential decoherence of the Werner parameter in memory and over channels, and analytic swapping rules. The model is designed for speed and scalability, but it does not include a built-in dynamic planner for per-hop protocol choice or multipartite GHZ and CV distillation.
SeQUeNCe [2] is an event-driven quantum network simulator that includes an entanglement management module with BBPSSW-style purification and entanglement swapping. Purification is available as a protocol component (BBPSSWProtocol and BBPSSW_BDS), and tutorials demonstrate fixed round schemes on short repeater chains. The Bell diagonal variant can be configured to behave as BBPSSW (with twirling) or DEJMPS (without twirling). Twirling applies random local Pauli rotations to symmetrize the state, effectively erasing specific error biases to yield a standard Werner state characterized solely by fidelity. SeQUeNCe exposes a method that computes the updated fidelity for a given input fidelity based on these configurations. However, SeQUeNCe does not provide a global path-level planner that jointly optimizes per-hop method, number of rounds, and end-to-end latency or resource cost.
@@
Beyond simulators, recent work on network-level purification and scheduling has studied how depolarizing noise, gate errors, storage times, and routing policies affect performance on complex topologies [18]. These works typically consider fixed purification protocols and focus on where to purify and how often. A complementary line of work optimizes repeater chain policies or cutoff thresholds using Markov decision processes or dynamic programming, but often without explicit purification choices. Y. Shi et al. [17] proposed a purification protocol selection module for bipartite links that uses density matrix simulations under realistic errors to build heuristics for choosing BBPSSW, DEJMPS, or other protocols at the link layer. Their module optimizes over a local configuration space per hardware instance but does not combine these choices into a path-level planner that accounts for swapping and end-to-end objectives.
@@ IV-B1 Static Recurrence
On each link $\ell$, APC represents the current pair by either a scalar fidelity $F$ and Werner parameter $w$ or by a Bell-diagonal vector $\bm{\lambda}$, depending on the configured physical model. Conversion between fidelity and Werner parameter is given by Eq. 13.
@@
for Werner inputs using (2)–(3), and, where configured, Bell-diagonal updates $(\bm{\lambda},p_{\text{succ}})\mapsto(\bm{\lambda}^{\prime},p_{\text{succ}}^{\prime})$ using the DEJMPS relations. For $r$ rounds on link $\ell$ under protocol $P$, APC computes the sequence $F^{(j+1)}=F_{P}^{\prime}(F^{(j)})$ and the multi-round success probability $P_{\text{succ}}^{(r)}$ from (9). The expected raw entanglement cost per surviving pair after $r$ rounds is $C_{\text{pairs}}^{(r)}$ in (10).
@@ IV-B2 Local Operations and Memory Noise
$\mathcal{E}_{2}(\rho)=(1-\lambda_{2q})\rho+\lambda_{2q}\frac{\mathbb{I}_{4}}{4},$ (17)
where $\lambda_{2q}\in[0,1]$ is the depolarizing mixing strength. In KOSMOS, the user-facing parameters $p_{1}$, $p_{2}$, and $p_{\mathrm{meas}}$ denote Pauli-twirled per-operation error probabilities (single-qubit, two-qubit, and local measurement bit-flip, respectively), i.e., with total probability $p$ a uniformly random non-identity Pauli is applied (on the corresponding $n$-qubit subsystem) and with probability 1p1-p the identity is applied.
The backend converts these primitive Pauli error rates into an equivalent depolarizing strength for
$\Delta_{\lambda}(\rho)=(1-\lambda)\rho+\lambda\frac{\mathbb{I}_{d}}{d},\qquad d=2^{n},$ (18)
using the standard mapping
$\lambda=\frac{d^{2}}{d^{2}-1}\,p=\frac{4^{n}}{4^{n}-1}\,p.$ (19)
For the 1-qubit and 2-qubit cases used throughout APC this yields
@@
where the measurement term folds a local bit-flip probability into an effective two-qubit depolarizing kick (at leading order, two local readouts contribute to a joint check/BSM).
Independent depolarizing channels are composed by multiplying reliabilities, $1-\lambda_{\mathrm{tot}}=\prod_{i}(1-\lambda_{i})^{n_{i}}$. Under this conversion, $\lambda$ reaches 1 at $p\leq(4^{n}-1)/4^{n}$; for larger p the formal equivalence would yield $\lambda>1$, so we clamp $\lambda\in[0,1]$ as a numerical guardrail (the regimes of interest satisfy $p\ll 1$ for realistic quantum repeaters).”
Assuming $n_{1q}$ single-qubit gates, $n_{2q}$ two-qubit gates, and $n_{\mathrm{meas}}$ measurements per recurrence round, the per-round reliability factor is
$r_{\text{round}}$ $=r_{1}^{n_{1q}}r_{2}^{n_{2q}}r_{m}^{n_{\mathrm{meas}}},$ (21)
$r_{1}$ $=1-\lambda_{1q},\qquad r_{2}=1-\lambda_{2q},\qquad r_{m}=1-\lambda_{\mathrm{meas}},$ (22)
where $\lambda_{1q},\lambda_{2q},\lambda_{\mathrm{meas}}\in[0,1]$ (and hence $r_{1},r_{2},r_{m}\in[0,1]$).
@@
    • Werner tracking: apply the effective two-qubit depolarizing channel directly in the fidelity domain,
      $F^{\prime}\;\leftarrow\;(1-\lambda_{\mathrm{tot}})\,F^{\prime}+\lambda_{\mathrm{tot}}\frac{1}{4},$ (23)
      and, when needed, convert to/from the Werner parameter $w$ using Eq. 13,
@@
APC applies this relaxation over the effective dwell time of the kept pair during each round and during swap operations (local gate time plus any classical wait), as provided by the planner time model.
@@ IV-C Multipartite GHZ Stabilizer Distillation
APC includes a GHZ post-processing stage that can be invoked after bipartite arms from a hub to multiple leaf nodes have been purified. In the current implementation, we use a planner-grade stabilizer-pass acceptance model inspired by stabilizer-measurement-based GHZ distillation for graph states [10, 1, 6]. This model is designed to be conservative and computationally cheap rather than a circuit-exact simulation.
A noisy $N$-qubit GHZ state is modeled using an isotropic GHZ-diagonal approximation:
$\rho_{\mathrm{GHZ}}=F_{\mathrm{GHZ}}|\mathrm{GHZ}_{N}\rangle\langle\mathrm{GHZ}_{N}|+(1-F_{\mathrm{GHZ}})\,\sigma_{\perp},$ (26)
@@
While the stabilizer group is generated by $Z_{i}Z_{N}$ and $X^{\otimes N}$, our implementation models one full stabilizer pass per round consisting of (N1)(N-1) weight-2 $Z_{i}Z_{N}$ checks and one global $X^{\otimes N}$ check, i.e., $m=N$ checks per pass. We account for the resource overhead by charging approximately one ancilla resource per check, i.e., $m$ ancilla resources per GHZ pass.
To enable millisecond-scale planning, the backend uses an aggregate acceptance model rather than an explicit circuit simulation. Let $F\equiv F_{\mathrm{GHZ}}$ be the current GHZ fidelity, and define the (clamped) effective per-check imperfection
$\varepsilon\;\triangleq\;\mathrm{clip}_{[0,1]}\!\Big((1-F_{\mathrm{anc}})+p_{\mathrm{meas}}^{\mathrm{GHZ}}\Big),$ (27)
which aggregates ancilla infidelity and measurement bit-flip error into a single conservative parameter.
We approximate the probability that the ideal (target) stabilizer-eigenvalue string passes all $m$ checks as
$p_{\mathrm{keep,good}}\;\approx\;(1-\varepsilon)^{m}.$ (28)
For the noise component, we assume GHZ-diagonal weight is uniformly distributed over the $2^{m}-1$ non-target stabilizer-eigenvalue strings, and each check outcome is flipped independently with probability $\varepsilon$. Under this uniform-noise approximation, the average false-accept probability is
$p_{\mathrm{keep,bad}}\;\approx\;\frac{1-(1-\varepsilon)^{m}}{2^{m}-1},$ (29)
which is continuous in $\varepsilon$ and vanishes as $\varepsilon\to 0$.
The resulting (effective) block success probability and posterior fidelity are
$p_{\mathrm{succ}}^{\mathrm{GHZ}}$ $\approx F\,p_{\mathrm{keep,good}}+(1-F)\,p_{\mathrm{keep,bad}},$ (30)
$F_{\mathrm{GHZ,out}}$ $\approx\frac{F\,p_{\mathrm{keep,good}}}{p_{\mathrm{succ}}^{\mathrm{GHZ}}}.$ (31)
This formulation allows APC to budget both latency (roughly $m$ LOCC check outcomes per pass) and ancilla overhead without tracking the full $2^{N}$-dimensional density matrix.